OpenStudy (anonymous):

Solve using the quadratic formula: 3x^2-5x-4=0

OpenStudy (mertsj):

1. Factor the left side. 2. Set each factor equal to 0 3. Solve each equation.

OpenStudy (anonymous):

So you have \[ax ^{2}+bx+c=0\] the Quadratic Formula is \[\frac{ -b \pm \sqrt{b2-4ac} }{ 2a }\], so a=3, b=-5,c=-4. Fill in and solve.

OpenStudy (anonymous):

\[\frac{ -(-5)\pm \sqrt{(-5)^{2}-4(3)(-4)} }{2(3)}\]

OpenStudy (anonymous):

\[(-5)\sqrt{25-48}/2(3)\]

OpenStudy (anonymous):

\[(-5)\pm \sqrt{25--48}/2(3)\]

OpenStudy (anonymous):

help

OpenStudy (anonymous):